OpenStudy (anonymous):

Find the sum of a finite geometric sequence from n = 1 to n = 5, using the expression −3(4)^(n − 1).

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@gswag98

OpenStudy (anonymous):

is this like distributing

OpenStudy (anonymous):

I am not sure, but I don't think so.

OpenStudy (anonymous):

@camerondoherty

OpenStudy (camerondoherty):

They give you an equation: f(x)=-3(4)^(n-1) So what you need to do is sub n for 1, 2, 3, 4, 5 and solve for each. The you will get the numbers in your sequence... Here's an example: For n=1 f(x)=-3(4)^(n-1) f(1)=-3(4)^(1-1) f(1)=-3(4)^0 f(1)=-12^0 f(1)=-1

OpenStudy (camerondoherty):

So you do that with all the different vales of n that they tell you to solve for

OpenStudy (camerondoherty):

but dont put f(x) put g(x) to show it is a geometric sequence

OpenStudy (anonymous):

Wow, smart and beautiful, thank you!

OpenStudy (camerondoherty):

Lol thanx c;

OpenStudy (anonymous):

@aum can u help me out on my question when you get the time

OpenStudy (aum):

−3 * (4)^(n − 1) ----- (1) Put n = 1, -3 * 4^0 = -3 * 1 = -3 To find the other numbers, you can put n = 2, 3, 4 and 5 into (1) and do the calculation. But an easier method would be to notice that every time n increases by 1 it is equivalent to multiplying by 4 because the base is 4 and n occurs in the exponent. So the second number is: first number * 4 = -3 * 4 = -12 Third number is: -12 * 4 = -48 Fourth number is: -48 * 4 = -192 Fifth number is: -192 * 4 = -768 Add: -3 + (-12) + (-48) + (-192) + (-768) = -(3 + 12 + 48 + 192 + 768) = ?